[Hallc_running] [New Logentry] RC Daily Update

munoz at jlab.org munoz at jlab.org
Fri Dec 7 10:25:02 EST 2018

Logentry Text:
- We finished running the setting with SHMS at 9.79 deg overnight with 40 uA beam
- Around 8:30AM today we moved the SHMS to 6.79 deg, where it will remain until the beam energy change on Monday morning.
- We will take data at 32, 24 and 16 uA, cycling between these currents (roughly 8h at each current setting).
- The missing reference time errors observed last night are not completely understood, but they are marginal and should not compromise the data. They happen in only some runs, and when they do, they affect way less than 1% of the events. Eric/Brad will make an opportunistic access to look into it whenever is possible.
- Plan is still to run beam until 6AM on Monday. We'll go into restricted access once Radcon has time to survey the Hall (9 or 10AM).

RC Meeting:
- There will be a short RC meeting today at 4:15pm, second floor counting house.

Notices in effect for shift crews:
- Shift leaders, please fill out the check list once per shift.
- Shift leaders, please do a quick cross-check of the BTA numbers against your shift summary.
- Check that the pre-scaler gui is active and hasn't frozen / doesn't need re-started.
- Be on the lookout for any errors when running the replay, particularly errors concerning missing reference times.
- Note the new beam positions for this energy (on the whiteboard / note on wall)
- NB If the HV trips for SHMS S1Y5B contact Simona M. '''DO NOT change the HV from the nominal value!'''


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3636008
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181207/5479b745/attachment.html>

More information about the Hallc_running mailing list